OpenStudy (anonymous):

Please help!! will medal- Tell whether the system has one solution,infinity more solutions, or no solution. 23. y=2x-1 3y=6x-5 25. 2x-2y=5 y=x-4 27. y-x=5 3y=3x+15 29. 2y=x-2 3y=3/2x-3

OpenStudy (anonymous):

23 and 27

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

What? @Coolguy23780

OpenStudy (anonymous):

there's more solution's

OpenStudy (anonymous):

what about the other to

OpenStudy (anonymous):

I DONT KNOW ALL I KNOW IS THAT IT PROB MIGHT BE 23 AND 27 SORRY FOR CAPS

OpenStudy (anonymous):

Thanks can you help with some other stuff too? @Coolguy23780

OpenStudy (anonymous):

Have to medal me like you said

OpenStudy (anonymous):

ok bring the questions

OpenStudy (anonymous):

Solve each system using substitution- 11. x+y=8 y=3x 13. x+3=y 3x+4y=7 15. y=-2+6 3y-x+3=0 @Coolguy23780

OpenStudy (anonymous):

11 x and y = 8,13 x and y i think = 0,15 y=2-x=-3+3=0 thats all i got

OpenStudy (anonymous):

Thanks! help with more?

OpenStudy (anonymous):

@Coolguy23780

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

solve using elimination- 7. 3x+3y=27 x-3y=-11 9.2x+4y=22 2x-2y=-8 11.5x-4=0 3x+y=24

OpenStudy (anonymous):

7. A bicycle store costs $2400 per month to operate. The store pays an average of $60 per bike. The average selling price of each bike is $120. How many bikes must the store sell each month to break even?

OpenStudy (anonymous):

be back bathroom

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Coolguy23780

OpenStudy (anonymous):

@Coolguy23780 can you please help